5x=10y
How would you graph them? Thank you

Answer by mananth(16946)

5x=10y
y= 5x/10
y = 1/2 *x
if x = 4, y=2 (4,2)
if x= 6, y = 3 (6,3)
if x=-4, y=-2 (-4,-2)
plot the points and join them to get the line